Fermi National Laboratory
 

Accelerator Update for July 24, 2006 - July 26, 2006

 

Monday July 24

The day shift began with Operations monitoring store 4851 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, Main Injector (MI), and Recycler operating normally, with the NuMI horn cooling work continuing, and with MiniBooNE taking beam.

Operations began preparing the Recycler at 10:34 AM to stash antiprotons. The stash ended around 11:25 AM.

Operations prepared Pbar for access at 11:34 AM.

At 2:20 PM, Operations reported that the Tevatron (TeV) sector A1 pond pump strainer was clogged. FESS personnel gathered equipment for replacing the strainer. CDF and D0 turned off their silicon detectors for the duration of the work.

The A1 strainer was successfully replaced. CDF and D0 had their silicon detector back on by 2:50 PM.

Pbar came out of access and resumed stacking at 6:07 PM.

Tuesday July 25

The midnight shift began with Operations monitoring store 4851, with Pbar stacking, and with MiniBooNE taking beam.

Operations terminated store 4851 at 1:02 AM.

At 1:13 AM, Operations reported that the controls system (sequencers, heartbeat, datalogging) was not working. Parameter pages and alarms were still working. Operators contacted experts.

Operations prepared D0 for access at 1:21 AM.

A controls expert reported at 2:01 AM that a new system had been released; he retreated to the previous system and this resolved everything.

Operations prepared CDF for access at 3:39 AM.

At 4:54 AM, Operations reported a power glitch. The Recycler and Pelletron stayed up, but nearly everything else went down. Operations contacted many experts.

At 5:03 AM, the Duty Electrician reported that feeder 49 had tripped off.

Operations reported at 5:09 AM that the D0 and CDF backup generators came on.

At 5:40 AM, the High Voltage Coordinator reported that feeder 49 had a ground fault.

MI was ramping by 6:58 AM.

A cryo system expert at CHL reported at 7:12 AM that the TeV cryo system was mostly recovered.

The High Voltage Coordinator reported at 7:20 AM that he would connect CDF to feeder 47 and D0 to feeder 45. This was completed by 7:30 AM.

Operations began preparing the TeV at 7:35 AM for access.

Operations reported at 7:49 AM that the Booster RF was back on.

The TeV access ended at 11:27 AM.

Operations reported at 3:11 PM that they had established beam in Linac, Booster, MI, and to MiniBooNE.

Operations ramped the TeV up to 980 GeV at 3:13 PM.

Pbar began stacking at 3:34 PM.

CDF experimenters completed their access at 3:50 PM. D0 came out of access about an hour later.

Operations put the TeV through a dry squeeze at 5:45 PM.

At 6:13 PM, the TeV suffered a sector F2 quench. Operators discovered that the F17 kicker had failed to fire.

Operations put the TeV through a dry squeeze at 7:41 PM.

Operations put the TeV through a wet squeeze at 9:23 PM.

Operations put the TeV into shot setup at 9:46 PM.

Operations established store 4857 at 11:43 PM with an initial luminosity of 142.33E30.

Wednesday July 26

The midnight shift began with Operations monitoring store 4857 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

The Recycler began preparing to stash at 12:57 AM. The stash ended and Pbar resumed stacking at 1:38 AM.

The Recycler began preparing to stash at 5:07 AM. The stash ended and Pbar resumed stacking at 5:39 AM.

NuMI experts reported at 7:17 AM that they were starting their chillers.

The Plans for Wednesday

The plans for today are to stash, stack, and store. NuMI will begin taking beam this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/26/2006    email Fermilab